OpenStudy (akonkel):

WILL MEDAL AND FAN :) A copy machine makes 60 copies per minute. A second copy machine makes 80 copies per minute. The second machine starts making copies 2 minutes after the first machine starts. Both machines stop making copies 8 minutes after the first machine started. Together, the 2 machines made how many copies?

OpenStudy (akonkel):

@AdriannaD2016 @Ahsan730 @bibby @beastieman21 @bkeith2698 @brittanydosey @Compassionate @CausticSyndicalist @cupcakerain @chappy0729 @DanJS @dareintheren @ddubay28 @dmw6 @eliassaab @Ellia201 @e2020.sucks @faariat @falloutboi @fathimarushda @ganeshie8 @Godlovesme @gmitsch @GreatMath @hailiehargis02 @heydolly__ @highschooler99 @inkyvoyd @i_love_bands26 @julian25 @jojobuck @jonnyboy101 @JordanTheGirl @Kainui @kittiwitti1 @kennybm @kingMAN @LiteNing1337 @Leyla_12 @lhunter @lopez09 @myininaya @Mashy @monsterfox1997 @MTALHAHASSAN2 @naturalpuff @Novel9012 @neharikam @notjusttheanswer @One098 @pooja195 @pootypie @PlanSlam @princesskole33 @quickstudent @Rea201 @Radelp @RockerChic8 @raerae252 @SolomonZelman @skyphos @shelbydancer @Sadworld @TuringTest @TQKMB @TheRaggedyDoctor @Taylor15 @UCLA104 @Werewolfprincess @Wxlfz @xapproachesinfinity @xo_kansasprincess_xo @Xaze @yomamabf @1armywifey @123333333333333333333 could any of u help me

OpenStudy (anonymous):

1240 Copies

OpenStudy (one098):

turn the tagging a notch down.

OpenStudy (akonkel):

These are the only options i have 480 600 680 720 960

OpenStudy (akonkel):

lol no kindin :)

OpenStudy (rea201):

The correct answer is 960.

OpenStudy (rea201):

60*8=480+(80*6)=960

OpenStudy (akonkel):

r u sure? a lot of people had just randomly told me an answer then i got it wrong after i meddled and fanned them :(

OpenStudy (akonkel):

thanx :P

OpenStudy (rea201):

I am positive. Your welcome :)